### v1

# Improved Constructions of Reed-Solomon Codes with Optimal Repair Bandwidth

Abstract
Maximum-distance-separable (MDS) codes are widely used in distributed storage, yet naive repair of a single erasure in an $[n,k]$ MDS code downloads the entire contents of $k$ nodes. Minimum Storage Regenerating (MSR) codes (Dimakis et al., 2010) minimize repair bandwidth by contacting $d>k$ helpers and downloading only a fraction of data from each. Guruswami and Wootters first proposed a linear repair scheme for Reed-Solomon (RS) codes, showing that they can be repaired with lower bandwidth than the naive approach. The existence of RS codes achieving the MSR point (RS-MSR codes) nevertheless remained open until the breakthrough construction of Tamo, Barg, and Ye, which yields RS-MSR codes with subpacketization $\ell=s\prod_{i=1}^{n}p_{i}$, where $p_{i}$ are distinct primes satisfying $p_{i}\equiv 1\pmod{s}$ and s=d+1ks=d+1-k.
In this paper, we present an improved construction of RS-MSR codes by eliminating the congruence condition $p_{i}\equiv 1\pmod{s}$. Consequently, our construction reduces the subpacketization by a multiplicative factor of $\phi(s)^{n}$ ( $\phi(\cdot)$ is Euler’s totient function) and broadens the range of feasible parameters for RS-MSR codes.
I Introduction
The exponential growth of data in distributed storage systems has necessitated the use of efficient erasure codes to ensure reliability. Maximum Distance Separable (MDS) codes, particularly Reed-Solomon (RS) codes, are preferred due to their optimal storage efficiency. However, a major drawback of traditional MDS codes is the high repair bandwidth required to recover a single failed node.
The concept of Regenerating Codes, introduced by Dimakis et al. [3], established the fundamental tradeoff between storage overhead and repair bandwidth. Minimum Storage Regenerating (MSR) codes achieve the cut-set bound for repair bandwidth while maintaining the MDS property.
Assume the data are encoded by a code $\mathcal{C}$ over a finite field $F$, where each code symbol is either a single element of $F$ or an $\ell$-dimensional vector over $F$. In the former case, $\mathcal{C}$ is called a scalar code; in the latter case, $\mathcal{C}$ is called an array code with subpacketization $\ell$. Most MSR constructions have focused on array codes, and many interesting results have been obtained, see [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. In contrast, constructing scalar MDS codes, particularly Reed–Solomon codes, that achieve the cut-set bound remained an open problem for several years. A landmark result by Tamo, Ye, and Barg [1] provided the first and, to date, only explicit construction of $(n,k)$ Reed–Solomon codes that are MSR codes, utilizing the linear repair scheme for scalar codes proposed by Guruswami and Wootters [2]. Their construction, referred to as RS-MSR codes, selects evaluation points from a field extension determined by a set of primes $p_{i}$ satisfying $p_{i}\equiv 1\pmod{s}$, where s=d+1ks=d+1-k and $d$ denotes the number of helper nodes participating in the repair of a failed node. The resulting subpacketization is $\ell=s\prod_{i=1}^{n}p_{i}$. A natural question is whether RS-MSR codes can achieve a smaller subpacketization. In this work, we present a follow-up to the construction in [1] to answer this question affirmatively.
Our main contribution is a basis-transformation technique composed of three parts: Euclidean Square Partition, Reshape, and Interference, which transforms a basis of the extension field into another basis that enables optimal repair of RS codes. Our results can be regarded as a generalization of [1] by relaxing the restriction on the choice of the field extension degrees $p_{i}$, namely, $p_{i}\equiv 1\pmod{s}$, thereby reducing the subpacketization by a factor of $\phi(s)^{n}$, where $\phi(\cdot)$ denotes Euler’s totient function.
The remainder of this paper is organized as follows. Section II provides the necessary preliminaries on RS codes and linear repair schemes. Section III presents our main results. Section IV concludes this paper.
II Preliminaries
Throughout this paper, we use the following notation unless stated otherwise.
  • For integers $a\leq b$, define $a,b]:=\{a,a+1,\dots,b\}$.
  • For a positive integer $n$, define $n]:=\{1,\dots,n\}$.
  • For any array $A$ of size $a\times b$, we index its entries by $(i,j)\in[0,a-1]\times[0,b-1]$, and denote by $A(i,j)$ the entry in the $i$-th row and $j$-th column.
II-A RS Codes and Repair Bandwidth
Let $E$ be a finite field. Let $\Omega=\{\omega_{1},\omega_{2},\ldots,\omega_{n}\}\subseteq E$ be a set of $n$ pairwise distinct elements, and let $v=(v_{1},\ldots,v_{n})\in(E^{*})^{n}$, where $E^{*}:=E\setminus\{0\}$. The generalized Reed–Solomon code $\operatorname{GRS}_{E}(n,k,\Omega,v)\subseteq E^{n}$ with evaluation set $\Omega$ and column multipliers $v$ is defined as the set
$\left\{\left(v_{1}f(\omega_{1}),\ldots,v_{n}f(\omega_{n})\right)\in E^{n}:\ f\in E[x],\ \deg f\leq k-1\right\}.$
If $v=(1,\ldots,1)$, then the GRS code is called a Reed–Solomon code and is denoted by $\operatorname{RS}_{E}(n,k,\Omega)$. It is well known (see [20]) that
$\left(\operatorname{RS}_{E}(n,k,\Omega)\right)^{\perp}=\operatorname{GRS}_{E}(n,n-k,\Omega,u),$
where $u_{i}=\prod_{j\neq i}(\omega_{i}-\omega_{j})^{-1}$ for $i=1,\ldots,n$.
Given an $(n,k)$ MDS array code $\mathcal{C}$ over a finite field $F$ with subpacketization $\ell$ (i.e., each node stores a vector $c_{i}\in F^{\ell}$). Suppose that the $i$-th node $c_{i}$ is erased, the repair bandwidth $N(\mathcal{C},i,\mathcal{R})$ of $c_{i}$ is defined as the total amount of information (measured in symbols over $F$) downloaded from a helper set $\mathcal{R}\subseteq[n]\setminus\{i\}$ with $|\mathcal{R}|=d$ to reconstruct $c_{i}$. The cut-set bound of [3] states that the repair bandwidth $N(\mathcal{C},i,\mathcal{R})$ satisfies
$N(\mathcal{C},i,\mathcal{R})\geq\frac{d\ell}{d-k+1}.$ (1)
We call $\mathcal{C}$ an MSR code if it achieves (1) with equality for all $i\in[n]$.
Assume that $E$ is an extension field of $F$ with $E:F]=\ell$, then we can regard an RS code over $E$ as an MDS array code over $F$ with subpacketization $\ell$. An RS code over $E$ is called an RS-MSR code if it is an MSR code when we regard it as an MDS array code over $F$.
II-B Linear Repair Schemes for scalar MDS codes
A key result in [2] is that the repair of a failed node $c_{i}$ can be performed by downloading the values of $\ell$ dual codewords at the helper nodes. Specifically, let $\mathcal{C}^{\perp}$ be the dual code of $\mathcal{C}$. If there exists a set of dual codewords $\{h_{1},\dots,h_{\ell}\}\subset\mathcal{C}^{\perp}$ such that their $i$-th coordinates span $E$ over $F$, then the node $c_{i}$ can be repaired.
Theorem 1 ([2]).
Let $\mathcal{C}\subset E^{n}$ be a scalar linear MDS code of length $n$. Let $F$ be a subfield of $E$ such that $E:F]=\ell$. For a given $i\in[n]$ the following statements are equivalent.
  1. (1)
    There is a linear repair scheme of the node $c_{i}$ over $F$ such that the repair bandwidth $N(\mathcal{C},i,\mathcal{R})\leq b$.
  2. (2)
    There is a subset of codewords $P_{i}\subseteq\mathcal{C}^{\perp}$ with size $|P_{i}|=\ell$ such that
    ${\rm dim}_{\mathbb{F}}(\{x_{i}:x\in P_{i}\})=\ell,$
    and
    $b\geq\sum_{j\in[n]\setminus\{i\}}{\rm dim}_{F}(\{x_{j}:x\in P_{i}\}).$
II-C Reed-Solomon codes as MSR codes
In this subsection, we introduce the main idea of the construction of RS-MSR codes in [1]. Precisely, given any positive integers $k<d\leq n-1$, the goal is to construct an $(n,k)$ RS code $\mathcal{C}$ achieving the cut-set bound for the repair of any single node from any $d$ helper nodes.
Let $\mathbb{F}_{p}$ be a finite field of prime order $p$. Denote s:=dk+1s:=d-k+1 and let $p_{1},\dots,p_{n}$ be $n$ distinct primes such that
$p_{i}\equiv 1\;\text{mod}\,s\;\;\text{~for all~}i=1,2,\dots,n.$ (2)
For $i=1,\dots,n$, let $\alpha_{i}$ be an element of an extension field of $\mathbb{F}_{p}$ such that $\mathbb{F}_{p}(\alpha_{i}):\mathbb{F}_{p}]=p_{i}$, and let
$\mathbb{F}:=\mathbb{F}_{p}(\alpha_{1},\dots,\alpha_{n}).$
Then it can be checked that $\mathbb{F}:\mathbb{F}_{p}]=\prod_{i=1}^{n}p_{i}$. Next, we define $n$ distinct subfields $F_{i}$ of the field $\mathbb{F}$ and one extension field $\mathbb{E}$ of $\mathbb{F}$.
  1. 1.
    For $i=1,\dots,n$, define $F_{i}=\mathbb{F}_{p}(\{\alpha_{j}:j\neq i\})$. Note that $\mathbb{F}=F_{i}(\alpha_{i})$ and $\mathbb{F}:F_{i}]=p_{i}$.
  2. 2.
    The field $\mathbb{E}$ is defined to be the degree-$s$ extension of the field $\mathbb{F}$, i.e., there exists an element $\beta\in\mathbb{E}$ of degree $s$ over $\mathbb{F}$ such that $\mathbb{E}=\mathbb{F}(\beta).$ We also have $\mathbb{E}:F_{i}]=sp_{i}$ for all $i$.
The construction of RS-MSR codes in [1] is as follows.
Theorem 2 ([1]).
Let $k,n,d$ be any positive integers such that $k<d<n.$ Let $\Omega=\{\alpha_{1},\dots,\alpha_{n}\}$, where $\alpha_{i},i=1,\dots,n$ is an element of degree $p_{i}$ over $\mathbb{F}_{p}$ and $p_{i}$ is the $i$-th smallest prime that satisfies (2). The code $\mathcal{C}:=\text{\rm RS}_{\mathbb{E}}(n,k,\Omega)$ achieves the cut-set bound for the repair of any single node from any $d$ helper nodes. In other words, $\mathcal{C}$ is an $(n,k)$ MSR code with repair degree $d$ and subpacketization $s\prod_{i=1}^{n}p_{i}$.
The proof of Theorem 2 relies on the following lemmas.
Lemma 1 ([21]).
Let $E$ be an extension field of finite field $F$ with $E:F]=\ell$. Assume $\{\zeta_{1},\dots,\zeta_{\ell}\}$ and $\{\mu_{1},\dots,\mu_{\ell}\}$ are dual basis of $E$ over $F$, i.e., ${\rm tr}(\zeta_{i}\mu_{j})=\delta_{ij}$, where ${\rm tr}(\cdot)$ denote the trace function from $E$ to $F$ and $\delta_{ij}$ is the Kronecker delta. Then for any $\gamma\in E$,
$\gamma=\sum_{i=1}^{\ell}{\rm tr}(\zeta_{i}\gamma)\mu_{i}.$
For a vector space $V$ over a field $F$ and a set of vectors $A=\left(a_{1},\ldots,a_{l}\right)\subset V$, let $\operatorname{Span}_{F}(A)=\left\{\sum_{i=1}^{l}\gamma_{i}a_{i},\gamma_{i}\in F\right\}$ be the span of $A$ over $F$.
Lemma 2 ([1]).
Let $\alpha_{i},\beta,F_{i}$ be as defined previously. Given $i\in[n]$, define the following vector spaces over $F_{i}$:
$S_{i}^{(1)}=\operatorname{Span}_{F_{i}}\big(\beta^{u}\alpha_{i}^{u+qs},u\in[0,s-1],q\in[0,\frac{p_{i}-1}{s}-1]\big)$
$S_{i}^{(2)}=\operatorname{Span}_{F_{i}}\Big(\sum_{t=0}^{s-1}\beta^{t}\alpha_{i}^{p_{i}-1}\Big)$
$S_{i}=S_{i}^{(1)}+S_{i}^{(2)}.$
Then
$\dim_{F_{i}}S_{i}=p_{i}\text{ and }S_{i}+S_{i}\alpha_{i}+\dots+S_{i}\alpha_{i}^{s-1}=\mathbb{E}.$ (3)
In what follows, we present a compressed version of the proof of Theorem 2 from [1] to highlight the role of Lemma 2 in determining the subpacketization.
Let $\mathcal{C}={\rm RS}_{\mathbb{E}}(n,k,\Omega)$, $\mathbb{F}$ be a subfield of $\mathbb{E}$ with $\mathbb{E}:\mathbb{F}]=\ell$. Let $c=(c_{1},c_{2},\dots,c_{n})$ be a codeword of $\mathcal{C}$, assume that $c_{i}$ is erased, it can be regenerated by the following procedure:
  • (1)
    By Lemma 2, there exists a vector space $S_{i}$ over the field $F_{i}$ such that
    $\dim_{F_{i}}S_{i}=p_{i}\text{ and }S_{i}+S_{i}\alpha_{i}+\dots+S_{i}\alpha_{i}^{s-1}=\mathbb{E},$
    let $e_{1},\dots,e_{p_{i}}$ be an arbitrary basis of the subspace $S_{i}$ over the field $F_{i}$.
  • (2)
    Choose $R_{i}\subset[n]\setminus\{i\}$ with $|R_{i}|=d$, where d=s+k1d=s+k-1.
  • (3)
    Let $h(x)=\prod_{j\in[n]\setminus(R_{i}\cup\{i\})}(x-\alpha_{j})$, then for $0\leq t\leq s-1$, $(v_{1}\alpha_{1}^{t}h(\alpha_{1}),\dots,v_{n}\alpha_{n}^{t}h(\alpha_{n}))\in C^{\perp},$ hence $\sum_{j=1}^{n}c_{j}v_{j}\alpha_{j}^{t}h(\alpha_{j})=0$
  • (4)
    Assume that $R_{i}=\{i_{1},\dots,i_{d}\}$, then
    $\sum_{j=1}^{d}c_{i_{j}}v_{i_{j}}\alpha_{i_{j}}^{t}h(\alpha_{i_{j}})=0.$
  • (5)
    Let ${\rm tr}_{i}:={\rm tr}_{\mathbb{E}/F_{i}}$ denote the trace function from $\mathbb{E}$ to $F_{i}$. By Lemma 1, the set
    $\left\{{\rm tr}_{i}(c_{i}e_{m}v_{i}\alpha_{i}^{t}h(\alpha_{i}))\right\}_{0\leq t\leq s-1,1\leq m\leq p_{i}}$
    can be used to recover $c_{i}$, as $\left\{e_{m}v_{i}\alpha_{i}^{t}h(\alpha_{i}))\right\}_{0\leq t\leq s-1,1\leq m\leq p_{i}}$ forms a basis of $\mathbb{E}$ over $F_{i}$.
  • (6)
    For each $0\leq t\leq s-1$ and $1\leq m\leq p_{i}$,
    ${\rm tr}_{i}(c_{i}e_{m}v_{i}\alpha_{i}^{t}h(\alpha_{i}))$ $=-\sum_{j=1}^{d}{\rm tr}_{i}(c_{i_{j}}e_{m}v_{i_{j}}\alpha_{i_{j}}^{t}h(\alpha_{i_{j}}))$
    $=-\sum_{j=1}^{d}\alpha_{i_{j}}^{t}h(\alpha_{i_{j}}){\rm tr}_{i}(c_{i_{j}}e_{m}v_{i_{j}})$
    Hence, from each helper node, for example, the $i_{1}$-th node in $R_{i}$, one only need to calculate $\{{\rm tr}_{i}(c_{i_{1}}v_{i_{1}}e_{m})\}_{1\leq m\leq p_{i}}$, i.e. $p_{i}$ symbols form $F_{i}$. Consequently, the total repair bandwidth is $dp_{i}=\frac{sp_{i}d}{d+1-k}$, meet the cut-set bound (1).
Remark 1.
From the compressed proof of Theorem 2, the key to optimal repair is to find a $\mathbb{E}$-subspace $S_{i}$ satisfying (3). Lemma 2 provides such subspaces under the congruence condition $p_{i}\equiv 1\pmod{s}$, which yields an RS-MSR code with subpacketization $\ell=s(\prod_{i=1}^{n}p_{i}).$ In the next section, we show that this congruence condition can be removed, which further reduces the subpacketization.
III Main results
In this section, we generalize Lemma 2 by constructing an explicit $S$ satisfying (3) without assuming $p_{i}\equiv 1\pmod{s}$, thereby obtaining an RS–MSR code with smaller subpacketization $\ell=s\prod_{i=1}^{n}p_{i}$ for any choice of distinct primes $p_{1},\ldots,p_{n}>s$.
Lemma 3.
Assume that $p$ and $s$ are positive integers such that $p>s$. Let $\beta$ be a generator of $\mathbb{K}$ over $\mathbb{E}$ with $\mathbb{K}:\mathbb{E}]=s$, and let $\alpha$ be a generator of $\mathbb{E}$ over $\mathbb{F}$ with $\mathbb{E}:\mathbb{F}]=p$. Then there exists an $\mathbb{F}$-subspace $S\subseteq\mathbb{K}$ with $\dim_{\mathbb{F}}(S)=p$ such that
$\mathbb{K}=\bigoplus_{u=0}^{s-1}\beta^{u}\mathbb{E}=\sum_{u=0}^{s-1}\alpha^{u}S.$
The main idea for finding the desired subspace $S$ is to view the problem as a change of $\mathbb{F}$-basis. Viewing $\mathbb{K}$ as an $\mathbb{F}$-vector space of dimension $sp$, one checks that
$\bigl\{\alpha^{i}(\alpha\beta)^{j}:i=0,\dots,p-1,\ j=0,\dots,s-1\bigr\}$
forms an $\mathbb{F}$-basis of $\mathbb{K}$. We arrange it as a $p\times s$ array $B$ (Figure 1). For $(i,j)\in[0,p-1]\times[0,s-1]$,
$B(i,j)=\alpha^{i}(\alpha\beta)^{j}.$ (4)
Thus, proving Lemma 3 reduces to the problem of finding another $\mathbb{F}$-basis that can be arranged as an $s\times p$ array (Figure 2) such that the $\mathbb{F}$-span of the first row equals $S$, and each subsequent row is obtained by multiplying the first row by a successive power of $\alpha$.
[…]
Figure 1: Original ($p\times s$)
$\rightarrow\,?$
[…]
Figure 2: Target ($s\times p$)
We achieve this basis change via a transformation procedure composed of: Euclidean Square Partition, Reshape, and Interference.
Euclidean Square Partition
Let $p>s$, and write the Euclidean algorithm as
$p$ $=a_{0}s+b_{1},$ (5)
$s$ $=a_{1}b_{1}+b_{2},$
$b_{1}$ $=a_{2}b_{2}+b_{3},$
$\ \vdots$
$b_{m-2}$ $=a_{m-1}b_{m-1}+b_{m},$
$b_{m-1}$ $=a_{m}b_{m},$
where $b_{m}=\gcd(p,s)$.
Given a $p\times s$ array $A$, we partition its entries into square blocks by partitioning its index set
$\mathcal{I}(p,s)\triangleq[0,p-1]\times[0,s-1]$
into disjoint squares via the following deterministic recursive procedure.
Definition 1 (Euclidean Square Partition).
Let $h,w$ be positive integers and let
$\mathcal{I}(h,w)\triangleq[0,h-1]\times[0,w-1].$
We define a partition of $\mathcal{I}(h,w)$ into disjoint squares via the following recursive procedure. Let $t:=\min\{h,w\}$. Tile $\mathcal{I}(h,w)$ by placing as many $t\times t$ squares as possible along the longer side (left-to-right if $w\geq h$, and top-to-bottom if $h>w$). This leaves a (possibly empty) residual rectangle, denoted by $\mathcal{I}(h^{\prime},w^{\prime})$; if it is nonempty, apply the same procedure recursively to $\mathcal{I}(h^{\prime},w^{\prime})$. The resulting collection of squares forms a disjoint partition of $\mathcal{I}(h,w)$, called the Euclidean square partition of $\mathcal{I}(h,w)$ and denoted by $\mathcal{P}(h,w)$. Each square in $\mathcal{P}(h,w)$ is called a Euclidean square.
This partition induces a partition of the array $A$ into square subarrays: for each $Q\in\mathcal{P}(h,w)$, let
$A[Q]\triangleq\bigl(A(i,j)\bigr)_{(i,j)\in Q}.$
With a slight abuse of notation, we use the same symbol $Q$ to refer to both the index square $Q\subseteq\mathcal{I}(h,w)$ and the associated square block $A[Q]$ whenever no confusion arises.
Applying the above procedure to $(h,w)=(p,s)$ yields $\mathcal{P}(p,s)$. Moreover, the side lengths of the squares in $\mathcal{P}(p,s)$ form the multiset
$\Big\{\underbrace{s,\dots,s}_{a_{0}},\underbrace{b_{1},\dots,b_{1}}_{a_{1}},\dots,\underbrace{b_{m},\dots,b_{m}}_{a_{m}}\Big\},$
as induced by the Euclidean chain (5).
Reshape
Given the Euclidean Square Partition $\mathcal{P}(p,s)$, we define a reshape map $\mathsf{Reshape}_{p,s}(\cdot)$ that converts a $p\times s$ array into an $s\times p$ array by (i) taking a overall transpose, and then (ii) performing a block-wise transpose (as specified in Definition 2).
Definition 2 (Reshape).
Let $B$ be a $p\times s$ array and let $B^{\top}$ be the $s\times p$ array defined by $B^{\top}(i,j)=B(j,i)$. For each Euclidean square $Q$ in $B$ with top-left corner index $(x,y)$ and side length $t$, let $Q^{\top}$ be the corresponding $t\times t$ square in $B^{\top}$ with top-left corner index $(y,x)$. Define $R=\mathsf{Reshape}_{p,s}(B)$ as the array obtained from $B^{\top}$ by transposing every square $Q^{\top}$, equivalently, for every such $Q$ and all $0\leq u,v\leq t-1$,
R(y+u,x+v)=B(x+u,y+v).R(y+u,\ x+v)=B(x+u,\ y+v). (6)
Interference
Firstly, we provide the definition of the interference operation.
Definition 3.
Let $m$ be a positive integer, and let $x,j_{1},j_{2},\dots,j_{m}$ be non-negative integers. Denote $J$ as the set of all distinct integers among $j_{1},j_{2},\dots,j_{m}$. For any symbols $\alpha,\beta$, consider the sequence $\{a_{i}\}_{i=1}^{m}$ defined by
$a_{i}=\alpha^{x+i}\beta^{j_{i}},\qquad i=1,\dots,m,$
i.e., the exponents of $\alpha$ are consecutive. We define the interference operator $\mathrm{Int}(\cdot)$ by mapping each term $a_{i}$ to
$\mathrm{Int}(a_{i}):=\overline{a}_{i}=\alpha^{x+i}\left(\sum_{j\in J}\beta^{j}\right)$
We call $\overline{a}_{i}$ the interference associated with $a_{i}$, and we call $\{\overline{a_{i}}\}_{i=1}^{m}$ the interference of the sequence $\{a_{i}\}_{i=1}^{m}$.
Example 1.
If $a_{1}=\alpha^{2}\beta^{2}$, $a_{2}=\alpha^{3}\beta^{4}$, and $a_{3}=\alpha^{4}\beta^{4}$, then the interference of the sequence $\{a_{1},a_{2},a_{3}\}$ is
$\left\{\alpha^{2}(\beta^{2}+\beta^{4}),\alpha^{3}(\beta^{2}+\beta^{4}),\alpha^{4}(\beta^{2}+\beta^{4})\right\}.$
Execute the interference operator $\mathrm{Int}(\cdot)$ column-wise on the reshaped array $R$, and denote the resulting $s\times p$ array by $\overline{R}:=\mathrm{Int}(R)$.
The interference operation in Definition 3 requires that, within each column, the exponents of $\alpha$ form a consecutive sequence. This property is guaranteed by the following lemma.
Lemma 4 (Consecutive $\alpha$-exponents after the reshape).
For every $(i,j)\in[0,s-1]\times[0,p-1]$, we have
$R(i,j)=\alpha^{i+j}\,\beta^{\tau(i,j)}$ (7)
for some integer $\tau(i,j)\in[0,s-1]$ depending on $(i,j)$. In particular, for each fixed column index $j$, the $\alpha$-exponents appearing in the $j$-th column of $R$ are exactly
j,j+1,,j+s1,j,\ j+1,\ $\dots$,\ j+s-1,
hence form a consecutive sequence of length $s$.
Proof.
By Definition 2, for any Euclidean square $Q\in\mathcal{P}(p,s)$ with top-left corner index $(x,y)$ and side length $t$,
$R(y+u,\ x+v)=B(x+u,\ y+v),\qquad 0\leq u,v\leq t-1.$
Substituting the expression of $B$ in (4) gives
$R(y+u,\ x+v)=\alpha^{x+u}(\alpha\beta)^{y+v}=\alpha^{x+u+y+v}\beta^{y+v}.$
Let $i=y+u$ and $j=x+v$. Then $x+u+y+v=i+j$, and hence
$R(i,j)=\alpha^{i+j}\beta^{y+v},$
which proves (7) with $\tau(i,j)=y+v\in[0,s-1]$. For fixed $j$, the exponent of $\alpha$ in $R(i,j)$ equals $i+j$, so as $i$ runs from 0 to s1s-1 we obtain the consecutive set $\{j,\dots,j+s-1\}$. ∎
Remark 2.
$\mathrm{Int}(\cdot)$ acts as the identity on the first $a_{0}s$ columns of $R$.
The entire transformation procedure is illustrated in Fig. 3 using a concrete example.
[…]
Figure 3: The transformation procedure for $(p,s)=(7,5)$.
Define
$S:=\operatorname{Span}_{\mathbb{F}}\{\overline{R}(0,t):\,0\leq t\leq p-1\}.$ (8)
We claim that this subspace is exactly the desired $S$ in Lemma 3. Before proving this claim, we first give the explicit form of $S$.
Since $\mathrm{Int}(\cdot)$ acts as the identity on the first $a_{0}s$ columns of $R$, we decompose $S=S^{(1)}+S^{(2)},$ where
$S^{(1)}$ $=\operatorname{Span}_{\mathbb{E}}\{\overline{R}(0,qs+u):q\in[0,a_{0}-1],u\in[0,s-1]\},$
$S^{(2)}$ $=\operatorname{Span}_{\mathbb{E}}\{\overline{R}(0,a_{0}s+j):0\leq j\leq b_{1}-1\}.$
It is easy to verify that for $0\leq q\leq a_{0}-1$ and $0\leq u\leq s-1$,
$\overline{R}(0,qs+u)=R(0,qs+u)=\alpha^{qs}(\alpha\beta)^{u}.$
Hence $S^{(1)}$ is determined, and it suffices to determine $S^{(2)}$.
For $0\leq j\leq b_{1}-1$, define
$\mathcal{R}_{j}:=\overline{R}(0,a_{0}s+j),\\$
$j^{(0)}:=j,\qquad j^{(2t)}:=j^{(2t-2)}-\Big\lfloor\frac{j^{(2t-2)}}{b_{2t}}\Big\rfloor b_{2t}\quad(t\geq 1).$
By Definition 3, determining $\mathcal{R}_{j}$ amounts to identifying all $\beta$-exponents that appear in the $(a_{0}s+j)$-th column of $R$. We give an explicit expression for $\mathcal{R}_{j}$ and defer its proof to Appendix A.
Proposition 1.
For each $j\in\{0,1,\dots,b_{1}-1\}$,
$\mathcal{R}_{j}$ $=\alpha^{a_{0}s+j}\Bigg(\sum_{\begin{subarray}{c}1\leq t\leq\lceil m/2\rceil\\ \forall k\in\{1,\dots,t-1\}:\ \left\lfloor\frac{j^{(2k-2)}}{b_{2k}}\right\rfloor=a_{2k}\end{subarray}}\ \sum_{q=0}^{a_{2t-1}-1}\beta^{\left(\sum_{k=1}^{t-1}a_{2k-1}b_{2k-1}\right)+q\,b_{2t-1}+j^{(2t-2)}}$
$\hskip 73.97733pt+\,\sum_{\begin{subarray}{c}1\leq t\leq\lfloor m/2\rfloor\\ \forall k\in\{1,\dots,t-1\}:\ \left\lfloor\frac{j^{(2k-2)}}{b_{2k}}\right\rfloor=a_{2k}\\ \left\lfloor\frac{j^{(2t-2)}}{b_{2t}}\right\rfloor<a_{2t}\end{subarray}}\beta^{\left(\sum_{k=1}^{t}a_{2k-1}b_{2k-1}\right)+j^{(2t)}}\Bigg),$ (9)
Corollary 1.
When $m=1$, from Proposition 1 we have that for $j\in[0,b_{1}-1]$, $\mathcal{R}_{j}=\alpha^{a_{0}s}(\alpha\beta)^{j}\sum_{t=0}^{a_{1}-1}\beta^{tb_{1}}.$
Proof.
We substitute $m=1$ into (1). Since $\lceil m/2\rceil=1$ and $\lfloor m/2\rfloor=0$, the second sum in (1) is empty, and the first sum contains only the term $t=1$. For $t=1$ we have $\sum_{k=1}^{t-1}a_{2k-1}b_{2k-1}=0$, hence the exponents contributed by the unique vertical term are
$tb_{1}+j,\qquad t=0,1,\dots,a_{1}-1.$
Therefore,
$\mathcal{R}_{j}=\alpha^{a_{0}s+j}\sum_{t=0}^{a_{1}-1}\beta^{tb_{1}+j}=\alpha^{a_{0}s}(\alpha\beta)^{j}\sum_{t=0}^{a_{1}-1}\beta^{tb_{1}}.$
Remark 3.
Lemma 2 corresponds to the special case of this corollary where $b_{1}=1$.
Corollary 2.
When $m=2$, from Proposition 1 we have that for $j\in[0,b_{1}-1]$,
$\mathcal{R}_{j}=\alpha^{a_{0}s}(\alpha\beta)^{j}\Bigg(\sum_{t=0}^{a_{1}-1}\beta^{tb_{1}}+\beta^{a_{1}b_{1}-\left\lfloor j/b_{2}\right\rfloor b_{2}}\Bigg).$
Proof.
We substitute $m=2$ into (1). Here $\lceil m/2\rceil=1$ and $\lfloor m/2\rfloor=1$, hence both sums in (1) contain only the term $t=1$.
The vertical contribution (the first sum with $t=1$) is identical to the case $m=1$:
$\alpha^{a_{0}s+j}\sum_{t=0}^{a_{1}-1}\beta^{tb_{1}+j}.$
The horizontal contribution: since $b_{1}=a_{2}b_{2}$ and $0\leq j\leq b_{1}-1$, we have $0\leq\lfloor j/b_{2}\rfloor\leq a_{2}-1$, so the condition $\big\lfloor j^{(0)}/b_{2}\big\rfloor<a_{2}$ holds for all such $j$. Therefore, the horizontal term contributes the exponent
$a_{1}b_{1}+j^{(2)}.$
Combining both contributions yields
$\mathcal{R}_{j}=\alpha^{a_{0}s+j}\Bigg(\sum_{t=0}^{a_{1}-1}\beta^{tb_{1}+j}+\beta^{a_{1}b_{1}+j^{(2)}}\Bigg)=\alpha^{a_{0}s}(\alpha\beta)^{j}\Bigg(\sum_{t=0}^{a_{1}-1}\beta^{tb_{1}}+\beta^{a_{1}b_{1}-\lfloor j/b_{2}\rfloor b_{2}}\Bigg),$
where we used $\beta^{a_{1}b_{1}+j^{(2)}}=\beta^{j}\beta^{a_{1}b_{1}-\lfloor j/b_{2}\rfloor b_{2}}$.
It is easy to verify that the example in Fig 3 coincides with this corollary.
Here we provide a sketch of proof of Lemma 3 and an illustrative example. The complete proof will be given in Appendix B.
The sketch of the proof of Lemma 3: from Remark 2, we only need to prove that the last $b_{1}$ columns of $R$ can be recovered by elements from vector space spanned from $\overline{R}$. Which can be done column by column and in each column the recovery order is from top to bottom. For example, in Fig. 3, the recovery proceeds as follows:
  1. 1.
    $R(0,5)=\overline{R}(0,5)-\overline{R}(3,2)-\overline{R}(1,4)$,
  2. 2.
    $R(1,5)=\overline{R}(1,5)-\overline{R}(4,2)-\overline{R}(2,4)$,
  3. 3.
    $R(2,5)=\overline{R}(2,5)-\alpha^{7}-\overline{R}(3,4)$,
  4. 4.
    $R(3,5)=\overline{R}(3,5)-\alpha^{8}-\overline{R}(4,4)$,
  5. 5.
    $R(4,5)=\overline{R}(4,5)-\alpha^{9}-\alpha^{9}\beta^{2}$,
  6. 6.
    $R(0,6)=\overline{R}(0,6)-\overline{R}(3,3)-\overline{R}(2,4)$,
  7. 7.
    $R(1,6)=\overline{R}(1,6)-\overline{R}(4,3)-\overline{R}(3,4)$,
  8. 8.
    $R(2,6)=\overline{R}(2,6)-\alpha^{8}\beta-\overline{R}(4,4)$,
  9. 9.
    $R(3,6)=\overline{R}(3,6)-\alpha^{9}\beta-R(4,5)$,
  10. 10.
    $R(4,6)=\overline{R}(4,6)-\alpha^{10}\beta-\alpha^{10}\beta^{3}$.
The above recovery is valid because
$\alpha^{7},\alpha^{8},\alpha^{9}$ $\in\operatorname{Span}_{\mathbb{E}}\{\{\overline{R}(t,0)\}_{t=0}^{4},R(0,5),R(1,5)\},$
$\alpha^{9}\beta^{2}$ $\in\operatorname{Span}_{\mathbb{E}}\{\{\overline{R}(t,2)\}_{t=0}^{4},R(2,5),R(3,5)\},$
$\alpha^{8}\beta,\alpha^{9}\beta,\alpha^{10}\beta$ $\in\operatorname{Span}_{\mathbb{E}}\{\{\overline{R}(t,1)\}_{t=0}^{4},R(0,6),R(1,6)\},$
$\alpha^{10}\beta^{3}$ $\in\operatorname{Span}_{\mathbb{E}}\{\{\overline{R}(t,3)\}_{t=0}^{4},R(2,6),R(3,6)\},$
We are now ready to present our improved construction of RS–MSR codes. The optimal repair scheme is essentially the same as that in [1], see Section II-C.
Theorem 3.
Let $k,n,d$ be positive integers such that $k<d<n$, and set s:=d+1ks:=d+1-k. Given a finite field $\mathbb{F}_{q}$, let $p_{i}$ be the $i$-th smallest prime greater than $s$. Choose elements $\alpha_{1},\dots,\alpha_{n}$ such that $\deg_{\mathbb{F}_{q}}(\alpha_{i})=p_{i}$ for $i=1,\dots,n$, and set $\mathbb{F}:=\mathbb{F}_{q}(\alpha_{1},\dots,\alpha_{n})$. Let $\mathbb{E}$ be an extension field of $\mathbb{F}$ such that $\mathbb{E}:\mathbb{F}]=s$, and let $\Omega=\{\alpha_{1},\dots,\alpha_{n}\}\subseteq\mathbb{F}$. Then the code $\mathcal{C}:=\mathrm{RS}_{\mathbb{E}}(n,k,\Omega)$ is an $(n,k)$ MSR code over $\mathbb{F}_{q}$ with repair degree $d$ and subpacketization $\ell=s\prod_{i=1}^{n}p_{i}$.
Remark 4 (The deduction of subpacketization).
Let $p_{i}$ denote the $i$-th prime greater than a constant $s$, and let $p_{i}^{\prime}$ denote the $i$-th prime satisfying $p_{i}^{\prime}\equiv 1\pmod{s}$. From the prime number theorem and the prime number theorem in arithmetic progressions [22], it can be shown that for $i\to\infty$,
$p_{i}\sim i\log i,~~p_{i}^{\prime}\sim\varphi(s)\,i\log i.$
Therefore,
$\frac{\ell_{\mathrm{TYB}}=\prod_{i=1}^{n}p_{i}^{\prime}}{\ell_{\mathrm{new}}=\prod_{i=1}^{n}p_{i}}\sim\varphi(s)^{n}.$
IV Conclusion
In this paper, we improve the construction of RS-MSR codes by showing that the previously imposed congruence condition on the primes $p_{i}$ is unnecessary, thereby reducing the subpacketization. A natural direction for future work is to determine whether the subpacketization can be reduced further, or to establish new lower bounds proving the tightness of our result.
Appendix A Proof of Proposition 1
Recall that determining $\mathcal{R}_{j}$ amounts to identifying the $\beta$-exponents appearing in the $(a_{0}s+j)$-th column of $R$, where $j\in\{0,1,\dots,b_{1}-1\}$. For simplicity, let $R_{0}$ denote the last $b_{1}$ columns of $R$. The following useful fact follows directly from the definition of the map $\mathsf{Reshape}_{p,s}(\cdot)$.
Fact: After the $\mathsf{Reshape}_{p,s}(\cdot)$, the squares in $R$ corresponding to Euclidean squares in $B$ form an Euclidean square partition of $R$, so does the squares in $R_{0}$, hence we can denote the set of all squares in $R_{0}$ by $\mathcal{P}(s,b_{1})$.
From the proof of Lemma 4, the exponents of $\beta$ inside each column of any Euclidean square $P$ in $R_{0}$ are the same. This motivates the following definition..
Definition 4 (Intersection indices $(y,v)$ for a fixed column).
Fix $j\in\{0,1,\dots,b_{1}-1\}$. Define $\mathcal{YV}(j)$ to be the set of all pairs $(y,v)$ for which there exist a Euclidean square $Q\in\mathcal{P}(s,b_{1})$ with upper-left corner $(y,x)$ and side length $t$ such that
$j\in\{x,x+1,\dots,x+t-1\}\quad\text{and}\quad v:=j-x\in\{0,1,\dots,t-1\}.$
Equivalently, $(y,v)\in\mathcal{YV}(j)$ if and only if the column indexed by $j$ in $\mathcal{I}(s,b_{1})$ intersects a Euclidean square whose upper-left row coordinate is $y$ with within-square column index $v$.
From Lemma 4, it follows immediately that
Proposition 2 (Reduction of $\mathcal{R}_{j}$ to $\mathcal{YV}(j)$).
For each $j\in\{0,1,\dots,b_{1}-1\}$, we have
$\mathcal{R}_{j}=\alpha^{a_{0}s+j}\sum_{(y,v)\in\mathcal{YV}(j)}\beta^{\,y+v}.$ (10)
Hence the core task is to characterize $\mathcal{YV}(j)$ for each $j$. This is purely a question about the Euclidean square partition $\mathcal{P}(s,b_{1})$. Set $b_{0}:=s$, the recursive construction of $\mathcal{P}(s,b_{1})$ begins with $b_{0}=a_{1}b_{1}+b_{2}$ and alternates between two types of decomposition steps:
  • the (2t1)(2t\!-\!1)-st vertical decomposition step, induced by
    $b_{2t-2}=a_{2t-1}b_{2t-1}+b_{2t},$ (11)
    which places $a_{2t-1}$ squares of side length $b_{2t-1}$ along the vertical direction;
  • the $(2t)$-th horizontal decomposition step, induced by
    $b_{2t-1}=a_{2t}b_{2t}+b_{2t+1},$ (12)
    which places $a_{2t}$ squares of side length $b_{2t}$ along the horizontal direction.
These steps are applied to the successive residual rectangles produced in the recursion.
Fix $j\in\{0,1,\dots,b_{1}-1\}$ and track the fixed global column $a_{0}s+j$ through the Euclidean recursion of $\mathcal{P}(s,b_{1})$. At any stage of the recursion, we describe the position of this fixed global column by its local column coordinate within the current residual rectangle. Vertical decomposition steps only cut the residual rectangle along the vertical direction and therefore preserve its horizontal interval, hence they do not change the local column coordinate. In contrast, at the $(2t)$-th horizontal decomposition step induced by $b_{2t-1}=a_{2t}b_{2t}+b_{2t+1}$, the current residual rectangle of width $b_{2t-1}$ is partitioned, from left to right, into $a_{2t}$ squares of side length $b_{2t}$ and one residual sub-rectangle of width $b_{2t+1}$. Accordingly, we define the local column coordinate recursively by
$j^{(0)}:=j,\qquad j^{(2t)}:=j^{(2t-2)}-\Big\lfloor\frac{j^{(2t-2)}}{b_{2t}}\Big\rfloor b_{2t}\quad(t\geq 1),$ (13)
so that $j^{(2t)}$ is the remainder of $j^{(2t-2)}$ upon division by $b_{2t}$.
Key observation: the column $a_{0}s+j$ enters the (2t1)(2t\!-\!1)-st vertical decomposition step if and only if, at each preceding horizontal decomposition step $(2k)$ with $k\in\{1,\dots,t-1\}$, the column lies in the residual sub-rectangle (rather than in one of the $a_{2k}$ squares of side length $b_{2k}$), i.e.,
$\forall k\in\{1,\dots,t-1\}:\ \Big\lfloor\frac{j^{(2k-2)}}{b_{2k}}\Big\rfloor=a_{2k}.$ (14)
Then we have:
  • (Vertical case: squares produced at the (2t1)(2t\!-\!1)-st vertical step). At the (2t1)(2t\!-\!1)-st vertical step, the recursion places $a_{2t-1}$ squares of side length $b_{2t-1}$ stacked vertically. Since these squares cover the full current width $b_{2t-1}$, the column intersects all of them, with the same within-square column index $v=j^{(2t-2)}$. Their upper-left row coordinates are
    $y=\Big(\sum_{k=1}^{t-1}a_{2k-1}b_{2k-1}\Big)+q\,b_{2t-1},\qquad q=0,1,\dots,a_{2t-1}-1.$
    Hence $\mathcal{YV}(j)$ contains the $a_{2t-1}$ pairs
    $(y,v)=\Bigg(\Big(\sum_{k=1}^{t-1}a_{2k-1}b_{2k-1}\Big)+q\,b_{2t-1},\ j^{(2t-2)}\Bigg),\qquad q=0,1,\dots,a_{2t-1}-1.$
  • (Horizontal terminating case: the unique square produced at the $(2t)$-th horizontal step). Now consider the subsequent $(2t)$-th horizontal step. If
    $\Big\lfloor\frac{j^{(2t-2)}}{b_{2t}}\Big\rfloor<a_{2t},$
    then the column lies in one of the $a_{2t}$ squares of side length $b_{2t}$ produced in this step, hence intersects exactly one such square. Its upper-left row coordinate equals
    $y=\sum_{k=1}^{t}a_{2k-1}b_{2k-1},$
    and the within-square column index is $v=j^{(2t)}$. Thus $\mathcal{YV}(j)$ contains the additional (terminating) pair
    $(y,v)=\Bigg(\sum_{k=1}^{t}a_{2k-1}b_{2k-1},\ j^{(2t)}\Bigg).$
    If instead $\big\lfloor j^{(2t-2)}/b_{2t}\big\rfloor=a_{2t}$, then the column lies in the residual sub-rectangle and this horizontal step contributes no pair, as the recursion continues to $(2t+1)$-th vertical decomposition.
Collecting the contributions from the vertical case (for all reachable depths $t$) and from the horizontal terminating case (whenever termination occurs at step $(2t)$) yields the following closed form:
$\mathcal{R}_{j}$ $=\alpha^{a_{0}s+j}\Bigg(\sum_{\begin{subarray}{c}1\leq t\leq\lceil m/2\rceil\\ \forall k\in\{1,\dots,t-1\}:\ \left\lfloor\frac{j^{(2k-2)}}{b_{2k}}\right\rfloor=a_{2k}\end{subarray}}\ \sum_{q=0}^{a_{2t-1}-1}\beta^{\left(\sum_{k=1}^{t-1}a_{2k-1}b_{2k-1}\right)+q\,b_{2t-1}+j^{(2t-2)}}$
$\hskip 73.97733pt+\,\sum_{\begin{subarray}{c}1\leq t\leq\lfloor m/2\rfloor\\ \forall k\in\{1,\dots,t-1\}:\ \left\lfloor\frac{j^{(2k-2)}}{b_{2k}}\right\rfloor=a_{2k}\\ \left\lfloor\frac{j^{(2t-2)}}{b_{2t}}\right\rfloor<a_{2t}\end{subarray}}\beta^{\left(\sum_{k=1}^{t}a_{2k-1}b_{2k-1}\right)+j^{(2t)}}\Bigg),$
which finished the proof of Proposition 1.
Appendix B Proof of Lemma 3
Let
$K=\sum_{u=0}^{s-1}\alpha^{u}S=\operatorname{Span}_{\mathbb{F}}\{\overline{R}(u,v):(u,v)\in\mathcal{I}(s,p)\}.$
From previous analysis, we only need to prove that for $(i,j)\in\mathcal{I}(s,b_{1})$, $R(i,a_{0}s+j)\in K$.
Fix $(i,j)\in\mathcal{I}(s,b_{1})$, let $\tau(\cdot,\cdot)$ be as in Lemma 4, then
$R(i,j):=\alpha^{a_{0}s+i+j}\beta^{\tau(i,a_{0}s+j)}.$
Equivalently, if $(i,j)$ lies in the unique Euclidean square whose upper-left corner is $(y,x)$, then Lemma 4 gives $\tau(i,a_{0}s+j)=y+(j-x)=j+(y-x)$, hence
$R(i,j)=\alpha^{a_{0}s+i}(\alpha\beta)^{j}\beta^{\,y-x}.$
Since $\mathcal{R}_{j}\in S$, for every $i\in[0,s-1]$ we have $\alpha^{i}\mathcal{R}_{j}\in K$. By Proposition 1 together with Lemma  4, the element $\alpha^{i}\mathcal{R}_{j}$ has the form
$\alpha^{i}\mathcal{R}_{j}=\alpha^{a_{0}s+i}(\alpha\beta)^{j}\Big(\beta^{y-x}+\sum_{\begin{subarray}{c}Q\text{ intersects col }j\\ Q\neq Q_{0}\end{subarray}}\beta^{y_{Q}-x_{Q}}\Big),$
where $Q_{0}$ is the Euclidean square containing $(i,j)$ and, for each other intersected square $Q$, $(y_{Q},x_{Q})$ denotes its upper-left corner. Consequently,
$R(i,j)=\alpha^{i}\mathcal{R}_{j}-\alpha^{a_{0}s+i}(\alpha\beta)^{j}\sum_{\begin{subarray}{c}Q\text{ intersects col }j\\ Q\neq Q_{0}\end{subarray}}\beta^{y_{Q}-x_{Q}}.$ (15)
Therefore it is enough to show that for every $Q\neq Q_{0}$ intersecting column $j$,
$\alpha^{a_{0}s+i}(\alpha\beta)^{j}\beta^{y_{Q}-x_{Q}}\in K.$ (16)
We prove (16) by induction on the column index $j$ ordered by the sequence of quotients appearing in the Euclidean recursion. Let $j^{(0)}:=j$ and $j^{(2t)}$ be defined by (13). Order columns by the lexicographic order of the finite sequence
$\Big(\Big\lfloor\frac{j^{(0)}}{b_{2}}\Big\rfloor,\ \Big\lfloor\frac{j^{(2)}}{b_{4}}\Big\rfloor,\ \Big\lfloor\frac{j^{(4)}}{b_{6}}\Big\rfloor,\ \ldots\Big),$
where the sequence stops when the reachability condition 14 first fails. Assume that for every column $j^{\prime}$ preceding $j$ in this order and every $i^{\prime}\in[0,s-1]$, we already have $R(i^{\prime},j^{\prime})\in K$.
Fix an intersected square $Q\neq Q_{0}$ in column $j$. If $Q$ is a terminating square arising from a horizontal decomposition step, then its upper-left column coordinate is of the form $x_{Q}\equiv 0\pmod{b_{2t}}$ for the corresponding step size $b_{2t}$, and the contribution of $Q$ in $\alpha^{i}\mathcal{R}_{j}$ is the term $\alpha^{a_{0}s+i}(\alpha\beta)^{j}\beta^{y_{Q}-x_{Q}}$ (this is exactly the terminating term in (1)). Choose an integer $z\geq 1$ such that $(i+zb_{2t},\,j-zb_{2t})\in\mathcal{I}(s,b_{1})$ lies in the same terminating square pattern at that step (such a $z$ exists because the terminating square has side length $b_{2t}$ and column $j$ lies in a block with $\lfloor j^{(2t-2)}/b_{2t}\rfloor>0$ at that step). Then, by the definition of $r(\cdot,\cdot)$ via Lemma 4 and the identity
$\alpha^{a_{0}s+i}(\alpha\beta)^{j}\beta^{y_{Q}-x_{Q}}=\alpha^{a_{0}s+i+zb_{2t}}(\alpha\beta)^{j-zb_{2t}}\beta^{y_{Q}-(x_{Q}-zb_{2t})}=\alpha^{a_{0}s+(i+zb_{2t})+(j-zb_{2t})}\beta^{\tau(i+zb_{2t},a_{0}s+j-zb_{2t})},$
we obtain
$\alpha^{a_{0}s+i}(\alpha\beta)^{j}\beta^{y_{Q}-x_{Q}}=R(i+zb_{2t},\,j-zb_{2t}).$
Moreover, $j-zb_{2t}$ precedes $j$ in the above lexicographic order because the quotient $\lfloor j^{(2t-2)}/b_{2t}\rfloor$ strictly decreases while all earlier quotients remain unchanged. Hence, by the induction hypothesis, $R(i+zb_{2t},\,j-zb_{2t})\in K$, which proves (16) for terminating squares.
If $Q$ is a vertical square in the Euclidean partition, then its contribution in (1) belongs to one of the vertical progressions at some vertical step, and the corresponding exponent difference $y_{Q}-x_{Q}$ differs from the target difference yxy-x by a positive multiple of the vertical step size at that stage. In this case, the element $\alpha^{a_{0}s+i}(\alpha\beta)^{j}\beta^{y_{Q}-x_{Q}}$ can be rewritten as
$\alpha^{a_{0}s+i}(\alpha\beta)^{j}\beta^{y_{Q}-x_{Q}}=(\alpha\beta)^{j+(y_{Q}-x_{Q})}\alpha^{a_{0}s+i-(y_{Q}-x_{Q})}.$
Since $Q\neq Q_{0}$ and $(i,j)\in Q_{0}$, the row index $i$ lies strictly above the vertical block indexed by $Q$ at that step, hence $0\leq a_{0}s+i-(y_{Q}-x_{Q})\leq a_{0}s-1$. Therefore $(\alpha\beta)^{j+(y_{Q}-x_{Q})}\alpha^{a_{0}s+i-(y_{Q}-x_{Q})}\in K$ by $\sum_{i=0}^{s-1}\alpha^{i}S^{(1)}\in K$, proving (16) for vertical squares.
Thus every term in the sum on the right-hand side of (15) belongs to $K$, so (15) implies $R(i,j)\in K$. By induction over the above column order, we have $R(i,j)\in K$ for all $(i,j)\in\mathcal{I}(s,b_{1})$, which finished the proof.